# Statistical Complexity of Heterogeneous Geometric Networks

### v2

@@ Abstract
Heterogeneity and geometry are key explanatory components underlying the structure of real-world networks. The relationship between these components and the statistical complexity of networks is not well understood. We introduce a parsimonious normalised measure of statistical complexity for networks. The measure is trivially 0 in regular graphs and we prove that this measure tends to 0 in Erdös-Rényi random graphs in the thermodynamic limit. We go on to demonstrate that greater complexity arises from the combination of hierarchical and geometric components to the network structure than either on their own. Further, the levels of complexity achieved are similar to those found in many real-world networks. We also find that real world networks establish connections in a way which increases complexity and which our null models and a range of attachment mechanisms fail to explain. This underlines the non-trivial nature of statistical complexity in real-world networks and provides foundations for the comparative analysis of network complexity within and across disciplines.
@@ I Introduction
When it comes to studying complexity in networks, we are concerned with the complexity of interactions– essentially, how diverse the connectivity patterns in the network are. While others borrow from the algorithmic view of complexity [6, 7], here we are concerned with the statistical complexity of networks. A notable early work on statistical complexity of networks introduced a measure called the network diversity score and provided a comprehensive overview of the problem [8]. Another work considered statistical complexity in networks from an information theoretic angle, multiplying Jensen Shannon divergence of a network with network entropy [9]. In this study, we establish a normalisation of Hierarchical Complexity (HC) as a network statistical complexity measure. In contrast to the network diversity score, HC is a parsimonious calculation of a single feature of a network which can be applied to individual networks and does not require a-priori categorisation of the complexity of the networks in question [10]. In contrast to the measure in [9] it does not require a reference graph, displays strong independence to network density and vanishes in the thermodynamic limit for Erdös-Rényi random graphs.
@@
A major issue with the generalisability of the HC measure is that it is not normalised by number of nodes or number of edges. Two similarly derived networks with different numbers of nodes or edges can be expected to have different values of HC. This paper addresses this issue by introducing a normalised HC measure. We show mathematically that this measure is bounded above by $2$ and satisfies the statistical complexity definition of being asymptotically zero for Erdös-Rényi random graphs (an appropriate equivalent to randomness in dynamical systems). We then go on to explore results of this normalisation on different types of random graphs with the two most evident structural properties relevant to real-world networks, hierarchy and geometry. Hierarchy here relates to the distribution of node fitness [16] or popularity [17]. Geometry relates to the latent space of similarities between nodes [18, 19]. Combining hierarchy and geometry successfully captures many of the properties of real-world networks [20], but whether these properties are enough to explain the statistical complexity of networks is not known. After this we go on to explore normalised HC in real-world networks. Here, an unexpected relationship between HC and density is noted. Finally, we explore explanations for this relationship by applying different kinds of attachment mechanisms based on degrees and overlap of node neighbourhoods. Combined, the results reveal non-trivial hierarchical complexity in real-world networks and we open the way for more reliable and robust applications of hierarchical complexity across network domains.
@@ II Theory
Key themes within this work are encapsulated within the image in Fig. 1 which illustrates hierarchical complexity arising from a combination of geometric and hierarchical structure. Henceforth, unless specified otherwise, let $G$ be a graph with $n$ nodes, $m$ edges and density $d$.
@@ II-B Normalised Hierarchical Complexity
Considering the minimal elements of each column and applying a similar argument to above we can also deduce that $2m^{\prime}\geq\alpha$, where $m^{\prime}=\frac{n(n-1)}{2}-m$ is the number of non-edges. Therefore, $\min(m,m^{\prime})\geq\alpha$. Moreover, $\max(m,m^{\prime})\geq\frac{n(n-1)}{4}$, since either at least half the edges are there or half are not.
@@ III-A Network models
III-B Attachment mechanisms
We used four different attachment mechanisms to observe the effect on hierarchical complexity of increasing density in different ways on real-world networks.
@@ III-B Attachment mechanisms
Random attachment: edge probabilities are uniform.
Hierarchical attachment: similar in fashion to preferential attachment for addition of new nodes to a network, the probability of a new edge is proportional to the sum of the degrees of the nodes:
@@
Similarity attachment: for $g_{i},g_{j}$ the neighbourhoods of nodes $i$ and $j$, the probability of edge $(i,j)$ occurring is proportional to the Jaccard index of their neighbourhoods:
@@
Combined hierarchical and similarity attachment: the probability of a new edge is proportional to the intersection of their neighbourhoods, essentially removing the size normalisation of the Jaccard index:
@@
so that the probability of connection is dependent on the size of the neighbourhoods and the overlap of the neighbourhoods.
To get the probabilities, $p_{ij}$, we divide each individual measurement (e.g., $k_{i}+k_{j}$ for hierarchical attachement) by the sum over all available measurements (e.g., $p_{ij}=(k_{i}+k_{j})/\sum_{(i,j)\notin\mathcal{E}}(k_{i}+k_{j})$). Edges are then randomly selected based on these probability spaces $\{p_{ij}\}_{(i,j)\notin\mathcal{E}}$.
@@ III-C Data
Data for replication in attachment mechanism experiment
For studying the attachment mechanisms, we used a replication dataset from the ICON corpus consisting of 139 networks mostly describing biological, social and technological phenomena [40, 15]. These ranged from $n=50$ to $n=3155$ with a mean of 341 $\pm$ 462. Densities ranged from $d=0.0011$ to $d=0.3884$ with a mean of 0.0578 $\pm$ 0.0717.
@@ IV-B Normalisation results on increasing number of nodes
We would expect networks of the same structure to have the same, or at least similar, HC. However, what do we mean by the “same structure”? If two networks have a different number of nodes they inherently have a different structure. In fact, the size of the network is related to the complexity, because as the size of random graphs increase their uniformity increases, this is due to the inverse relationship between variance and the sample mean. As such, we would expect that the normalised HC will decrease slightly as the number of nodes increases, but two sufficiently large networks of similar structure, but different size, would have similar complexity values.
@@ IV-C Effect of degree heterogeneity on hierarchical complexity
Here we study the change in complexity as we increase heterogeneity among the RHGGs. We generated 1000 RHGGs of size $n=1000$ and $d\sim U[0,1]$. The heterogeneity was randomly determined in each iteration with $\sigma_{h}\sim U[0,1]$. Results are shown in Fig 4.
Figure 4: Top, average normalised hierarchical complexity per degree of 1000 heterogeneous geometric graphs with with $n=1000$, $d\sim U[0,1]$, $\sigma_{h}\sim U[0,1]$. Bottom, global normalised hierarchical complexity plotted against heterogeneity of these networks.
The results show that normalised HC has a consistent range of values across most degrees for heterogeneous geometric graphs, tapering off quickly at either end towards 1 and $n$. It is beneficial that the measure does not have a positive or negative relationship with degree so that the measure does not overly emphasise any particular range of degrees. Further, the fact the measure quickly tapers off towards very small degrees protects the measure from being influenced by the high levels of uncertainty of sampling at these small sample sizes.
HC is generally strongest in models with $\sigma_{h}$ between 0.2 and 0.4 and decreases towards low and high heterogeneity. When $\sigma_{h}$ is high the network becomes dominated by the hierarchical relationships, which should make the network more ordered. On the other hand, with low $\sigma_{h}$ the model gets closer to a random geometric graph which has low values of HC as shown in Fig 2. This is consistent with our expectations of HC being a statistical measure of complexity in networks.
@@ IV-D Hierarchical complexity in large real-world networks
We compared unnormalised and normalised HC in the twenty large networks detailed in Table I. The results are shown in Table II. The values of the unnormalised HC range from 4 in the power grid network to 87858 in the German twitch network, a difference by a factor of nearly 22000. This range highlights the lack of utility of an unnormalised and unbounded measure for comparisons of different networks. On the other hand, the values of normalised HC range from 0.0022 in the power grid network up to 0.0830 in the Portugese twitch network, a difference by a factor of just under 38.
We can see that the normalised measure allows us to compare between these networks more clearly. For example, consider the twitch DE, twitch FR and twitch RU networks, which have the largest unnormalised HC values. These three networks are all constructed in the same way (friendships in twitch) and have similar densities (0.0034,0.0053, and 0.0039, respectively), and yet the twitch DE unnormalised HC value is 6 times bigger than the twitch RU unnormalised HC, suggesting that the twitch DE network is significantly more complex than the twitch RU network. However, when normalised these three networks all have very similar HC values, suggesting they have very similar levels of complexity, as one would expect.
Next consider the other three twitch networks: ES, PTBR and ENGBE. We can see that twitch ES and twitch ENGBE have very similar unnormalised HC values, but when normalised the twitch PTBR is twice as complex as twitch ES. This is not surprising considering they have markedly different densities, and we will see in the next section that even after normalisation the network density is correlated to the HC value in real world networks. We also see that the twitch ENGBE has much lower complexity than the other twitch networks, which is likely explained by the lower density of the network.
So we can see that our normalisation reduces the scale of the difference between complexities of networks, and allow us to better compare networks of different sizes. However, we also see that the density of the network still correlates with the normalised HC value.
@@ IV-E Growth of hierarchical complexity in real world networks
We found that hierarchical complexity was positively correlated with network density in the twenty networks from Table I ($\rho=0.7203$, $p=0.0005$), Fig. 5, top-left. We confirmed this observation with a network dataset of 139 smaller networks ($n\in[50,3155]$, $\rho=0.5325$, $p=1.5\times 10^{-11}$) [40, 15], Fig. 5, top-right. To discount the potential confounding effect of network size, $n$, on these correlations, we implemented linear regression on network density with network size as a predictor and found that the correlations of the residuals of the regression with HC were still significant in both cases– $\rho=0.6526$, $p=0.0023$ for the 20 large networks and $\rho=0.2294$, $p=0.0066$ for the 139 small-to-medium sized networks.
@@
Figure 5: Top row shows scatterplots visualising the positive association between density and hierarchical complexity in real-world networks. Spearman’s correlation coefficient and associated $p$-value shown inset. Bottom row shows average results of the values of HC as we increase density of networks according to the attachment mechanisms as described in the legend.
@@
At the same time, we have shown that hierarchical complexity shows strong normalisation with respect to density for many types of graph. The relationship between density and hierarchical complexity in real world networks is therefore unlikely due to a lack of normalisation, but is a true relationship requiring a mechanistic explanation. To try to explain this relationship we applied several attachment algorithms, as described in section III.B, to real-world networks to artificially increase their density and see if any would consistently lead to the targeted increase in HC.
Starting from the original networks, we increased density according to the four different attachment mechanisms in Section III.B from 0.1% up to a 2% increase in density in steps of 0.1%. Normalised HC was recorded at each step. The average results across the network datasets are shown in Fig. 5, bottom row. Patterns replicate in both datasets. Most evidently, none of the attachment mechanisms could explain an increase in complexity with increasing density as seen across the real-world networks. If anything, all mechanisms showed a decrease in complexity as density was increased.
Notably, the hierarchical & similar attachment mechanism which promotes connections primarily among those with most connections and those with greatest overlap of connections, did exhibit the least decrease in both datasets, as may be expected from our results in the RHGG models, but it still was not able to explain an increase in complexity.

### v3

@@ Abstract
Degree heterogeneity and latent geometry, also referred to as popularity and similarity, are key explanatory components underlying the structure of real-world networks. The relationship between these components and the statistical complexity of networks is not well understood. We introduce a parsimonious normalised measure of statistical complexity for networks. The measure is trivially 0 in regular graphs and we prove that this measure tends to 0 in Erdös-Rényi random graphs in the thermodynamic limit. We go on to demonstrate that greater complexity arises from the combination of heterogeneous and geometric components to the network structure than either on their own. Further, the levels of complexity achieved are similar to those found in many real-world networks. However, we also find that real-world networks establish connections in a way which increases complexity and which our null models fail to explain. We study this using ten link growth mechanisms and find that only one mechanism successfully and consistently replicates this phenomenon– probabilities proportional to the exponential of the number of common neighbours between two nodes. Common neighbours is a mechanism which implicitly accounts for degree heterogeneity and latent geometry. This explains how a simple mechanism facilitates the growth of statistical complexity in real-world networks.
Author Summary
A statistically complex system is one which is neither regular nor random, but contains diversity in components and structure. This departs from algorithmic complexity which describes how difficult it is to explain information, but which is maximal for uniformly random information. We provide a definition of statistical complexity for networks and propose a normalised measure which satisfies that definition. We go on to explore the relationship between statistical complexity and the two major components thought to underlie network structure– the popularity of nodes in making connections (degree heterogeneity) and the geometric similarity of nodes. We find that the statistical complexity of real-world networks agrees with a model which combines both components. We then notice a positive relationship between the density of links in real-world networks and their statistical complexity, which is not present in our modelling. We find that we can replicate this relationship, however, by growing the network using link probability which is based on a pair of nodes’ number of common neighbours. We conclude that statistical complexity is a natural by-product of uncomplicated network mechanisms, returning to the old adage that complexity arises from simplicity.
@@ I Introduction
When it comes to studying complexity in networks, we are concerned with the complexity of interactions– essentially, how diverse the connectivity patterns in the network are. While others borrow from the algorithmic view of complexity [6, 7], here we are concerned with the statistical complexity of networks. A notable early work on statistical complexity of networks introduced a measure called the network diversity score and provided a comprehensive overview of other complexity and entropy measures of networks and their limitations[8]. Another work considered statistical complexity in networks from an information theoretic angle, multiplying Jensen Shannon divergence of a network with network entropy [9]. Neither of these works, however, provides a treatment of statistical complexity as previously described, and the measurements have limitations partly owing to lack of normalisation to network size and/or density. In this study, we establish a normalisation of Hierarchical Complexity (NHC) as a network statistical complexity measure. In contrast to the network diversity score, NHC is a parsimonious calculation of a single feature of a network, rather than the product of four different measures [10]. In contrast to the measure in [9] it does not require a reference graph, displays strong independence to network density and vanishes in the thermodynamic limit for Erdös-Rényi random graphs.
@@
A major issue with the generalisability of the HC measure is that it is not normalised by number of nodes or number of edges. Two similarly derived networks with different numbers of nodes or edges can be expected to have different values of HC. This paper addresses this issue by introducing a normalised HC (NHC) measure. We show mathematically that this measure is bounded above by $2$ and satisfies the statistical complexity definition of being asymptotically zero for Erdös-Rényi random graphs (an appropriate equivalent to randomness in dynamical systems). We then go on to explore results of this normalisation on different types of random graphs with the two most evident structural properties relevant to real-world networks, hierarchy and geometry. Hierarchy here relates to the distribution of node fitness [16] or popularity [17]. Geometry relates to the latent space of similarities between nodes [18, 19]. Combining hierarchy and geometry successfully captures many of the properties of real-world networks [20], but whether these properties are enough to explain the statistical complexity of networks is not known. After this we go on to explore NHC in real-world networks. Here, an unexpected relationship between NHC and density is noted. Finally, we explore explanations for this relationship by applying different kinds of link growth mechanisms based on degrees and overlap of node neighbourhoods. Combined, the results reveal non-trivial hierarchical complexity in real-world networks and we open the way for more reliable and robust applications of hierarchical complexity across network domains.
@@ II Theory
Key themes within this work are encapsulated within the image in Fig. 1 which illustrates hierarchical complexity arising from a combination of geometric and hierarchical structure. Henceforth, unless specified otherwise, let $G$ be a graph with $n$ nodes, $m$ edges and density $d=2m/n(n-1)$.
@@ II-B Normalised Hierarchical Complexity
Considering the minimal elements of each column and applying a similar argument to above we can also deduce that $2m^{\prime}\geq\alpha$, where $m^{\prime}=\frac{n(n-1)}{2}-m$ is the number of non-edges. Therefore, $\min(m,m^{\prime})\geq\frac{\alpha}{2}$. Moreover, $\max(m,m^{\prime})\geq\frac{n(n-1)}{4}$, since either at least half the edges are there or half are not.
@@ III-A Network models
III-B Link growth mechanisms
We used ten different link growth mechanisms to observe the effect on NHC of increasing density in different ways on real-world networks.
@@ III-B Link growth mechanisms
Random growth
edge probabilities are uniform.
Popularity growth
similar in fashion to preferential attachment for addition of new nodes to a network, the probability of a new edge is proportional to the sum of the degrees of the nodes:
@@
Similarity growth
for $g_{i},g_{j}$ the neighbourhoods of nodes $i$ and $j$, the probability of edge $(i,j)$ occurring is proportional to the Jaccard index of their neighbourhoods:
@@
Popularity $\times$ similarity growth
the probability of a new edge is proportional to the intersection of their neighbourhoods, essentially removing the size normalisation of the Jaccard index:
@@
so that the probability of connection is dependent on the size of the neighbourhoods and the overlap of the neighbourhoods. Note, popularity and similarity growth is what is commonly known as the common neighbours algorithm, as it just counts the number of shared neighbours two nodes have to predict whether they will become connected in the future [31, 32]. While simple at face value, we can clearly see from the above that this works well to account for both popularity and similarity components of a network.
For each of the three latter approaches (popularity, similarity, and popularity $\times$ similarity) we took three different approaches to deciding on links: random probabilistic selection, random exponentiated probabilistic selection, and deterministic rank-based selection.
Probabilistic
To get the probabilities, $p_{ij}$, we divide each individual measurement (e.g., $k_{i}+k_{j}$ for hierarchical attachement) by the sum over all available measurements for example,
$p_{ij}=(k_{i}+k_{j})/\sum_{(i,j)\notin\mathcal{E}}(k_{i}+k_{j}).$ (17)
Edges are then randomly selected based on these probability spaces $\{p_{ij}\}_{(i,j)\notin\mathcal{E}}$.
Exponentiated probabilistic
This approach is taken to better differentiate between the strong and weak potential links in the probability space and so make it more likely for stronger potential links to be selected. Here, we simply take the exponentials of the probabilities before normalisation: $p_{ij}=\exp(k_{i}+k_{j})/\sum_{(i,j)\notin\mathcal{E}}\exp(k_{i}+k_{j})$.
Deterministic
Here, we simply take the top $x$ probabilities as new links. This is typically how link prediction would be done.
These three approaches span from the more randomised, to the more rigid, with exponentiated probabilistic growth taking the middle ground.
@@ III-C Data
Data for replication in link growth mechanism experiment
For studying the link growth mechanisms, we used a replication dataset from the ICON corpus consisting of 139 networks mostly describing biological, social and technological phenomena [42, 15]. These ranged from $n=50$ to $n=3155$ with a mean of 341 $\pm$ 462. Densities ranged from $d=0.0011$ to $d=0.3884$ with a mean of 0.0578 $\pm$ 0.0717.
@@ IV-A Normalisation results on random models
In the Supplementary Information section SI1, we explore the suitability of two relevant previously proposed network measures for assessing statistical complexity of networks using similar analyses as above. We find that neither meet the requirements expected of being 0 for regular and ER random graphs. We also find that they are not useful for distinguishing between different types of models which should be distinguishable in terms of statistical complexity.
@@ IV-B Normalisation results on increasing number of nodes
We would expect networks of the same structure to have the same, or at least similar, NHC. However, what do we mean by the “same structure”? If two networks have a different number of nodes they inherently have a different structure. In fact, the size of the network is related to the complexity, because as the size of random graphs increase their uniformity increases, this is due to the inverse relationship between variance and the sample mean. As such, we would expect that the NHC will decrease slightly as the number of nodes increases, but two sufficiently large networks of similar structure, but different size, would have similar complexity values.
@@ IV-C Effect of degree heterogeneity on hierarchical complexity
Here we study the change in complexity as we increase heterogeneity among the RHGGs. We generated RHGGs of size $n=1000$ and $d\sim U[0,1]$. The heterogeneity was determined with 100 realisations of the model for each $\sigma_{h}=0.01,0.02,\dots,2$. Results are shown in Fig 4.
Figure 4: Top, average normalised hierarchical complexity per degree of heterogeneous geometric graphs with with $n=1000$, $d\sim U[0,1]$, and $\sigma_{h}=0.01,0.02,\dots,2$. One hundred realisations are created for each $\sigma$. Bottom, global normalised hierarchical complexity plotted against heterogeneity of these networks.
The results show that NHC has a consistent range of values across most degrees for heterogeneous geometric graphs, tapering off quickly at either end towards 1 and $n$. It is beneficial that the measure does not have a positive or negative relationship with degree so that the measure does not overly emphasise any particular range of degrees. Further, the fact the measure quickly tapers off towards very small degrees protects the measure from being influenced by the high levels of uncertainty of sampling at these small sample sizes.
NHC is generally strongest in models with $\sigma_{h}$ between 0.2 and 0.4 and decreases towards low and high heterogeneity. When $\sigma_{h}$ is high the network becomes dominated by the hierarchical relationships, which should make the network more ordered. On the other hand, with low $\sigma_{h}$ the model gets closer to a random geometric graph which has low values of NHC as shown in Fig 2. This is consistent with our expectations of NHC being a statistical measure of complexity in networks.
In section SI2 of the Supplementary Information, we apply NHC to the non-uniform Popularity Similarity Optimisation model [43] to see how it fairs on another model which directly utilises degree heterogeneity and latent geometry. We find similar patterns as for RHGGs, that the model achieves highest NHC in a middle ground of degree heterogeneity. We also explore the effect of varying the clustering in the network and the number of communities through additional parameters of the model. We find that decreasing the clustering by randomising more connections decreases NHC, and that introduction strong community blocks in the geometry also decreases NHC. These are broadly consistent with our expectations of a statistical complexity metric.
@@ IV-D Hierarchical complexity in large real-world networks
We compared unnormalised and NHC in the twenty large networks detailed in Table I. The results are shown in Table II. The values of the non-normalised HC range from 4 in the power grid network to 87858 in the German twitch network, a difference by a factor of nearly 22000. This range highlights the lack of utility of an unnormalised and unbounded measure for comparisons of different networks. On the other hand, the values of NHC range from 0.0022 in the power grid network up to 0.0830 in the Portugese twitch network, a difference by a factor of just under 38.
We can see that the normalised measure allows us to compare between these networks more clearly. For example, consider the twitch DE, twitch FR and twitch RU networks, which have the largest non-normalised HC values. These three networks are all constructed in the same way (friendships in twitch) and have similar densities (0.0034,0.0053, and 0.0039, respectively), and yet the twitch DE non-normalised HC value is 6 times bigger than the twitch RU non-normalised HC, suggesting that the twitch DE network is significantly more complex than the twitch RU network. However, when normalised these three networks all have very similar NHC values, suggesting they have very similar levels of complexity, as one would expect.
Next consider the other three twitch networks: ES, PTBR and ENGBE. We can see that twitch ES and twitch ENGBE have very similar unnormalised HC values, but when normalised the twitch PTBR is twice as complex as twitch ES. This is not surprising considering they have markedly different densities, and we will see in the next section that even after normalisation the network density is correlated to the NHC value in real world networks. We also see that the twitch ENGBE has much lower complexity than the other twitch networks, which is likely explained by the lower density of the network.
So we can see that our normalisation reduces the scale of the difference between complexities of networks, and allow us to better compare networks of different sizes. However, we also see that the density of the network still correlates with the NHC value.
@@ IV-E Growth of hierarchical complexity in real world networks
We found that hierarchical complexity was positively correlated with network density in the twenty networks from Table I ($\rho=0.7203$, $p=0.0005$), Fig. 5, left. We confirmed this observation with a network dataset of 139 smaller networks ($n\in[50,3155]$, $\rho=0.5325$, $p=1.5\times 10^{-11}$) [42, 15], Fig. 5, right. To discount the potential confounding effect of network size, $n$, on these correlations, we implemented linear regression on network density with network size as a predictor and found that the correlations of the residuals of the regression with NHC were still significant in both cases– $\rho=0.6526$, $p=0.0023$ for the 20 large networks and $\rho=0.2294$, $p=0.0066$ for the 139 small-to-medium sized networks.
@@
Figure 5: Scatterplots visualising the positive association between density and hierarchical complexity in real-world networks. Spearman’s correlation coefficient and associated $p$-value shown inset. Bottom row shows average results of the values of NHC as we increase density of networks according to the link growth mechanisms as described in the legend.
@@
At the same time, we have shown that NHC shows strong normalisation with respect to density for many types of graph. The relationship between density and NHC in real world networks is therefore unlikely due to a lack of normalisation, but is a true relationship requiring a mechanistic explanation. To try to explain this relationship we applied ten link growth algorithms, as described in section III.B, to real-world networks to artificially increase their density and see if any would consistently lead to the targeted increase in NHC.
Firstly, we implemented link growth on 100 HRGGs of size $n=250$, density $d=0.05$ and density $\sigma=0.2$. Each link growth algorithm in Section III.B was applied for 50 iterations with 50 new links established each iteration. For the real-world networks, we applied the link growth algorithms to all of the 139 ICON networks and 17 of the 20 SNAP networks (omitting the 3 largest due to computational limitations). For these we used 10 iterations with 100 new links established each iteration. Results are shown in Fig. 6.
Figure 6: Average results of the values of NHC as we increase density of networks according to the link growth mechanisms as described in the legend. P stands for Popularity, S for Similarity and P $\times$ S for Popularity $\times$ Similarity. For the Heterogeneous geometric graphs we take the median over 100 iterations. For the ICON and SNAP networks we take the median of the z-scores of trajectories to avoid bias of networks with high NHC.
There are two clear observations to be made from these results. Firstly, the only kind of link growth which provided increased NHC were those based on the common neighbours (here referred to as Popularity $\times$ Similarity) algorithm. This shows a clear favoured strategy for link growth in complex networks.
All of the random, Popularity only and Similarity only algorithms worked to decrease NHC and in similar amounts as the baseline uniformly random algorithm. There was no particular observable consistent difference between the Popularity and Similarity only mechanisms. However, the deterministic mechanisms typically did worse most of the time. Infact, in the HRGGs the Popularity only deterministic mechanism did much worse than the uniformly random approach. Anytime the algorithm showed a greater decrease than the uniformly random mechanism we can assume this is caused by a more ordered structure being enforced on the network, since the only NHC smaller than ER random graphs are regular graphs and highly organised graphs where all nodes of the same degree have the same or similar neighbourhood degree sequences [15].
Secondly, the only algorithm which consistently increased the NHC across the three datasets was the exponentiated probabilistic Popularity $\times$ Similarity algorithm. This algorithm takes the middle ground between the more random standard probabilistic algorithm and the deterministic link ranking algorithm. The probabilistic mechanism increased for the SNAP networks over 10 iterations, but decreased slightly in the ICON networks and moreso in the RHGGS with a similar trajectory as for uniformly random growth. The deterministic mechanism intially increased alot for the RHGG model before going into a steep decline, while it increased in the SNAP networks but decreased in line with the uniformly random mechanism in the ICON networks.